OpenStudy (ally):

how do I solve x^2/5-x^2+2/2=x^2/10-11?

OpenStudy (ally):

X^2/5-x^2-2/2=x^2/10-11

OpenStudy (anonymous):

you sure that you got the equation right?

OpenStudy (ally):

X^2/5-x^2+2/2=x^2/10-11

OpenStudy (anonymous):

\[x ^{2}/5 - x ^{2}/2 + 2/2 = x ^{2}/10 -11\] is this the question?

OpenStudy (anonymous):

couldnt make sense of 2/2 or the right side of the equation

OpenStudy (anonymous):

or are those supposed to be rational exponents?

OpenStudy (anonymous):

but that is weird too...come to think

OpenStudy (ally):

anfar has it right

OpenStudy (anonymous):

if that right x^2 = -40/3 which doesnt make sense if youre not using imaginary values

OpenStudy (anonymous):

x=5 or -5

OpenStudy (anonymous):

sory.the answer should be x= 6 or -6

OpenStudy (anonymous):

ally. in the right side, is it + 2/2 or -2/2